# IFT-UAM/CSIC-24-89 Insights into Dark Matter Direct Detection Experiments: Decision Trees versus Deep Learning

### v1

@@ 5.2 Exclusion limits
We computed the exclusion limits employing two different methods detailed in Appendix B that combine ML and traditional statistical procedures. The process can be summarized as
Instead of dealing with the original high-dimensional space, we work with the output of the binary classifiers described in previous sections, which is always one-dimensional.
We employ two statistical methods, a binned and an unbinned analysis. In the first one, we divide the output of the ML in a histogram and estimate the expected number of signal and background events in each bin. In the second method, we estimate the signal and background probability density functions with a non-parametric algorithm. To do so, in both methods, we feed the already trained classifier with only signal events and then with only background events.
To set exclusion limits, we want to assess the compatibility of the signal-plus-background hypothesis when the true data has only background events. Therefore, we prepare a set of samples, or pseudo-experiments, each one containing the expected total number of background events, .
@@
We fix 222In the test statistic formula, we use the expected signal rate that corresponds to the fixed DM mass value. Additionally, we employ a ML classifier that was trained to distinguish background vs signal events generated with the mentioned fixed DM mass. Later we will see that we can design a single ML algorithm that can be applied to a range of DM masses. and compute the significance for different values of . We set the exclusion limits at 90% confidence level for the cross-section value with .
@@
In Fig. 6 we show the exclusion limits for some examples considering XGBoost as classifier. The solid lines represent the expected 90 C.L. upper limit on , while the dashed curves the 1 sensitivity bands taking into account only statistical uncertainties. We also present as a gray band the expected limits set by XENON1T 1 ton yr [53]. On the left panel, we show the results for each representation without including the feature , for the binned analysis. Although there are several differences between the analysis applied in this work and the one performed by the XENON collaboration (inclusion of all systematic uncertainties, data-driven analysis of the backgrounds, statistical treatment done on the original cS1-cS2 plane instead of the ML output of different representations), we can see that our results are compatible with XENON1T which validates the procedure and the simulator as an accurate tool to describe the behaviour of the detector. As expected from the AUC values from Table 2, the cS1-cS2 representation provides the strongest constraints, while the raw data (images) the worst performance with weaker constraints by a factor . The rest of exclusion limits considering other representations or ML algorithms lie between those curves. Importantly, if we include the information about , the exclusion limits of every representation are increasingly similar, following the AUC results of Table 2.
@@
Regarding the impact of employing a binned or an unbinned statistical treatment, on the right panel of Fig. 6 we show a comparison between the procedures considering cS1-cS2. The unbinned method out-performs the binned one by a factor , and we have checked that this fact holds for every representation.
@@
We would like to highlight that we have applied our procedure to an ideal scenario, i.e. computing the exclusion limits for a fixed DM mass using a specific ML classifier trained with a signal dataset generated with the same WIMP mass that we want to test. We call this approach dependent. Since in a real experiment we do not know the true underlying model, we have trained a new ML classifier designed to distinguish between background events and a signal dataset generated with different WIMP masses, GeV (equal number of events per mass). Then, we call this approach independent. Although the ML classifier trained with a set of masses is not optimal compared with a ML classifier trained for a specific mass in the sense that the AUC decreases when tested with a new set of signal events with that specific mass, the independent approach would be the one useful in a real experiment since it is suited to be applied to a range of masses. In Fig. 7 we show a comparison between both approaches, left panel for binned analysis and right panel for the unbinned one considering cS1-cS2. We can see that although not optimal, the independent algorithm presents minimal differences with the dependent method. This is related to the excellent discrimination power of the classifiers that are able to learn the distribution of the backgrounds, and to the similarities of the signal when only the WIMP mass is changed.
@@ 6 Conclusions
In this study, we performed an extensive comparison between deep learning models, such as state-of-the-art transformers, traditional methods like Multilayer Perceptrons and Convolutional Neural Networks, and boosted decision models like XGBoost to improve the detection capabilities of liquid xenon time projection chamber experiments. We evaluated different data representations and found that simplified feature representations, especially corrected S1 and S2 signals (cS1 and cS2), are effective in obtaining important information for classification tasks.
Our results show that while complex transformer models offer promising performance, simpler models such as XGBoost can achieve comparable results when optimal data representations are used. This emphasizes the importance of feature engineering and data representation in machine learning applications for DM detection.
@@ Appendix A Signal rates
Appendix B Statistical treatment
@@
The likelihood ratio close to 1 implies good agreement between the data and the hypothesized value of , otherwise the ratio is close to 0. Then , with higher values of, corresponding to a greater incompatibility between data and .
@@
We can estimate numerically the test statistic with a finite dataset . The probability distribution of is going to be conditioned to the true signal strength that defines the mentioned dataset. Since to set exclusion limits we want to assess the compatibility of signal-plus-background hypothesis when the true data has only background events, we construct a dataset containing only background events, i.e. . In practice, we generate a large number of pseudo-experiments 333For exclusion, a pseudo-experiment is a set of background-only events such that its number of events is taken from a Poisson distribution with mean equal to the expected number of background events for the considered experiment. taken from the dataset , and compute for each one the value of of Eq. 8. With that procedure, we can obtain numerically the distribution of and the median expected exclusion significance is defined as

### v2

@@ 5.2 Exclusion limits
We computed the exclusion limits employing two different methods detailed in Appendix C that combine ML and traditional statistical procedures. The process can be summarized as
Instead of dealing with the data representation original high-dimensional space, we work with the output of the binary classifiers described in previous sections, which is always one-dimensional.
We employ two statistical methods, a binned and an unbinned analysis. In the first one, we divide the output of the ML in a histogram and estimate the expected number of signal and background events in each bin. In the second method, we estimate the signal and background probability density functions with a non-parametric algorithm. To do so, in both methods, we feed the already trained classifier with only signal events and then with only background events taken from an independent dataset, the test set.
To set exclusion limits, we want to assess the compatibility of the signal-plus-background hypothesis when the true data has only background events. Therefore, we prepare a set of samples, or pseudo-experiments, each one containing the expected total number of background events, , taken from the test dataset.
@@
We fix 111In the test statistic formula, we use the expected signal rate that corresponds to the fixed DM mass value. Additionally, we employ a ML classifier that was trained to distinguish background vs signal events generated with the mentioned fixed DM mass. Later we will see that we can design a single ML algorithm that can be applied to a range of DM masses. and compute the significance for different values of . We set the exclusion limits at 90% confidence level for the cross-section value with .
@@
It is crucial to highlight that since the statistic treatment to determine the exclusion limits is applied to the output of the ML binary classifier, the exclusion reach is determined by the performance of the mentioned classifier. This implies that among the 32 models presented in Table 2, we expect that the one with the highest AUC will yield the most stringiest limits. Additionally, we also expect that models with similar AUCs will yield similar constraints. As discussed in the previous section, the XGBoost method offers several advantages, including simplicity, speed and greater stability. Therefore, we present the exclusion limit analysis using only XGBoost classifiers. For a comparison of some examples using transformers see Appendix D, which illustrates that the choice of ML architecture results in negligible differences in the exclusion limit curves when the AUCs are similar.
@@
On the left panel of Fig. 6 we show the exclusion limits for some examples considering XGBoost as classifier. The solid lines represent the expected 90 C.L. upper limit on , while the dotted curves the 1 sensitivity bands taking into account only statistical uncertainties. We also present as a gray band the expected limits set by XENON1T 1 ton yr [54]. Although there are several differences between the analysis applied in this work and the one performed by the XENON collaboration (inclusion of all systematic uncertainties, data-driven analysis of the backgrounds, and most importantly the fact that the statistical treatment is carried out on the original cS1-cS2 plane instead of on the ML output of different representations), we can see that our results are compatible with XENON1T which validates the procedure and the simulator as an accurate tool to describe the behaviour of the detector. As anticipated, the model with one of the highest AUC values in Table 2, dataset D (AUC=0.980), provides the strongest constraints, while dataset A1 (AUC=0.919) the worst limits by a factor . We have checked that the rest of exclusion curves, considering other representations and ML algorithms, lie between those curves following the AUC order shown in Table 2.
Next we will focus on the dataset C3. This is one of the simplest representations with only 5 variables (4 main resumed features and , the time between the pulses), that still provides one of the highest AUC and therefore a high exclusion limit reach. On the other hand, dataset D yields similar results with even less features (cS1 and cS2), and is the dataset usually employed by the experimental collaborations. However, it requires the determination of the correction factors to the original S1 and S2 signals, which implies a detailed detector characterization. This issue is absent with the dataset C3 case, as its features can be extracted directly from the raw signals. For completeness, in Appendix D we show the results using the well-known dataset D and compare them with the dataset C3. As expected, no significant differences are found.
@@
Regarding the impact of employing a binned or an unbinned statistical treatment, on the right panel of Fig. 6 we show a comparison between the procedures using dataset C3. The unbinned method out-performs the binned one by a factor , and we have checked that this fact holds for every representation. Nonetheless, it should be noted that the unbinned method is significantly more computationally expensive compared to the binned case (it takes about one to two orders of magnitude longer to compute).
@@
Finally, we would like to highlight that so far we have presented our procedure to an ideal scenario, i.e. computing the exclusion limits for a fixed DM mass using a specific ML classifier trained with a signal dataset generated with the same WIMP mass that we want to test. We call this approach dependent. Since in a real experiment we do not know the true underlying DM model, we have trained a new ML classifier designed to distinguish between background events and a signal dataset generated with different WIMP masses, GeV (equal number of events per mass). Then, we call this approach independent. Although a ML classifier trained with a set of masses may not be optimal compared with a ML classifier trained for a specific mass, the independent approach would be the one useful in a real experiment since it is suited to be applied to a range of masses. In Fig. 7 we show a comparison between both approaches, left panel for (the most efficient) binned analysis and right panel for (the best performing) unbinned one. We can see that for this particular physical problem, although not optimal, the independent algorithm presents minimal differences with the dependent method. This is related to both the excellent discrimination power of the classifiers that are able to learn the distribution of the backgrounds, and to the similarities of the signal when only the WIMP mass is changed.
@@ 6 Conclusions
In this study, we performed an extensive comparison between deep learning models, such as state-of-the-art transformers, traditional methods like Multilayer Perceptrons and Convolutional Neural Networks, and boosted decision models like XGBoost to improve the detection capabilities of liquid xenon time projection chamber experiments. We evaluated different data representations and found that simplified feature representations, especially dataset D (corrected S1 and S2 signals), and dataset C3 (4 main resumed features , the time difference between S1 and S2), are effective in obtaining important information for classification tasks. However, the latter representation avoids the need to determine the correction factors obtained through detailed detector characterization, and can be used as a simple cross-check.
Our results show that while complex transformer models offer promising performance, simpler and significantly faster models such as XGBoost can achieve comparable results when optimal data representations are used. This emphasizes the importance of feature engineering and data representation in machine learning applications for DM detection.
@@ Appendix A Signal rates
Appendix B Machine Learning training details
In this appendix we provide more details on the architecture and training of the ML classifiers. Since we studied 32 independently tuned data-algorithm combinations (not including several tests and combinations not shown in Table 2) in this section we present the details of some representative cases. For further information about the architectures, algorithm and dataset, the code used in this work is available at [60].
In Fig. 8 we show the train and validation loss as a function of the epoch during training, for the data representation dataset C3. To avoid overfitting an early stopping of 50 iterations was used in every ML algorithm to end the training procedure if the validation performance does not improve for the last given amount of epochs. Additionally, the command restorebestweights was set to True to save and use the weights from the best epoch during training (the one with the best performance), instead of using the weights from the last epoch that might be sub-optimal if training continued past the ideal point.
The XGBoost architecture has 12 hyperparameters, that we have tuned using a grid search. For the shown example in Fig. 8, the most relevant ones for our classification task are a maximum depth of a tree equal to 5, learning rate of 0.01, 2500 estimators, and a L2 regularization term on weights of 0.01, although we found that the default hyperparameter values provide near optimal results. As learning objective we choose binary:logistic and logloss as evaluation metric.
For the DNN example, we used 6 hidden layers with 128, 64, 32, 16, 8, and 4 units per layer, respectively, all with Leaky-ReLU as activation function with alpha of 0.05. These parameters, as well as the tuning of the activation function, have a significant impact in the network performance. Additionally, we set a dropout value of 0.2 after each layer, used a sigmoid function as activation in the final layer, defined binarycrossentropy as loss function for the training, using Adam as optimizer with its default values, and a batch size of 128. The number of training epoch was set high enough such that the early stopping algorithm finishes the fitting procedure. This results is a classifier with a total of 11809 trainable parameters.
The optimal CNN found for this example shares some hyperparameters with the mentioned DNN: the same optimizer, batch size, dropout values, and the same Leaky-ReLU activation function. We used 2 1D convolutional layers (since in this case we are using tabulated data) with 32 filters, and a kernel size of 3 and 2 for the first and second convolutional layer, respectively, a stride equal to 1, no padding, and an L2 kernel regularizer with value 0.001. After the convolutional layers we set a 1D max-pooling layer of size equal to 2, a dropout of 0.2, a flatten layer, and a fully connected layer of 32 units and Leaky-ReLU as activation function with alpha of 0.05, and dropout of 0.2. Finally, a sigmoid function as activation in the final layer was employed. Then, a network with a total of 3297 trainable parameters was generated.
The tabTransformer hyperparameters are set as follows: the embedding dimension (inputembeddim) is fixed at 128, which defines the size of the input embeddings for the features, providing a balance between model complexity and representational capacity. The number of attention heads (numheads) is set to 16, enabling the model to attend to different parts of the input representation simultaneously. Lastly, with numattnblocks=12, the model uses 12 stacked attention layers to capture hierarchical relationships between features while maintaining reasonable model depth.
To determine the dataset size needed, we performed a systematic study comparing the results using datasets of different length. For each ML classifier shown in Table 2, we computed the AUC 5 times per dataset length considered. We report the mean and one standard deviation. In Fig. 9 we show the AUC as a function of the train dataset size (which includes both background and signal events) for a few representative examples, noting that the rest of the ML classifiers follow the same tendency. As we can see, the AUC results are stable for datasets with events per class.
Appendix C Statistical treatment
@@
The likelihood ratio close to 1 implies good agreement between the data and the hypothesized value of , otherwise the ratio is close to 0. Then , with higher values of corresponding to a greater incompatibility between data and .
@@
We can estimate numerically the test statistic with a finite dataset . The probability distribution of is going to be conditioned to the true signal strength that defines the mentioned dataset. Since to set exclusion limits we want to assess the compatibility of signal-plus-background hypothesis when the true data has only background events, we construct a dataset containing only background events, i.e. . In practice, we generate a large number of pseudo-experiments 222For exclusion, a pseudo-experiment is a set of background-only events such that its number of events is taken from a Poisson distribution with mean equal to the expected number of background events for the considered experiment. taken from the dataset , and compute for each one the value of of Eq. 8. With that procedure, we can obtain numerically the distribution of and the median expected exclusion significance is defined as
@@
Appendix D Comparison of exclusion limits for several methods
In this appendix we compare the results for different data representation, ML algorithm and statistical treatment combinations. The exclusion limits shown on the first column of Fig. 10 are computed using the binned method, while the second column presents the results with the unbinned procedure. The first row corresponds to the dependent approach, where one specific ML classifier for each is trained, while the second row shows the independent strategy, using a single ML classifier and a single signal dataset that combines all the sets.
In particular, we compare the results using the data representation dataset C3 with XGBoost and tabTransformers, and dataset D with XGBoost. The first data representation is the simplest one since it is composed of 5 features that require only basic operations to be extracted from the detector response. Additionally this representation is among the best performing ones. For this case, we compare the simplest and fastest XGBoost algorithm against an state-of-the-art transformer. On the other hand, the second data representation is constructed with the well-known cS1, cS2 variables employed in the experimental collaborations analyses. It is composed of only 2 features, but requires a thorough detector characterization to determine the correction factors to the original S1 and S2 signals. As we can see from the figure, the simplest data-ML architecture combination provides the same exclusion power as the more complex ones.